The measure of the largest angle in a triangle is 20º less than the sum of the measures of the two smaller angles. The sum of the measures of the two larger angles is 90º greater than the measure of the smaller angle. Find the measure of each angle
Answer by mananth(16949)

let one small angle be x
& other be y
Largest angle = (x+y)-20
x+y+x+y-20=180
2x+2y=200
x+y=100
-------------
x+(x+y)-20=90+y
x+x+y-20=90+y
2x=110
x=55
y=45
the largest angle = 55+45-20=80
m.ananth@hotmail.ca